OpenStudy (anonymous):

Write two equations which represent the following two sentences: The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x.

OpenStudy (anonymous):

46=3x+4 x=14 x+x+y=46 28+y=46 y=18

OpenStudy (anonymous):

Okay thnk u @cxs